OpenStudy (anonymous):

how to graph the slope 2x -5y=20

OpenStudy (candyme):

2x-5y=20 First, you need to get the y by itself. -5y=-2x+20 5y=2x-20 Y= (2x-20)/5 Y=2/5x-4 To graph it you would put a point at (0,-4) then go up 2, right 5 from there.

OpenStudy (candyme):

See here

OpenStudy (candyme):

Your slope is 2/5